OpenStudy (lncognlto):

the iterative formula (which will be inserted) gives a sequence which converges to (formula 2).

OpenStudy (lncognlto):

\[x_{n+1} = \frac{ 1 }{ 2 } (\ln 9 - \ln x _{n})\]

OpenStudy (lncognlto):

formula 2: \[x = 9e ^{-2x}\]

OpenStudy (amistre64):

\[\frac{ 1 }{ 2 } (\ln 9 - \ln x _{n})\] \[\frac{ 1 }{ 2 }\ln \frac{9}{x_n}\] \[\ln \frac{3}{\sqrt{x_n}}\]

OpenStudy (lncognlto):

Ok, I got it, thanks!